OpenStudy (anonymous):

simplify the following expression: 8√8x^12

OpenStudy (anonymous):

x^12 under root becomes x^6. So the equation would be [x^6*8*\sqrt{8}\]

OpenStudy (anonymous):

You can also bring a 2 out of the square root because 2 * 4 = 8. Square root of 4 is 2.

OpenStudy (anonymous):

\[8\sqrt{8x^{12}}=8(\sqrt8)x^6=16\sqrt2 x^6\]. Since \(\sqrt{8}=2\sqrt{2}\).

OpenStudy (anonymous):

How did the 12 go to 6? Also where do I click to say that this questions has a good answer?

OpenStudy (anonymous):

Nevermind, found it! But how did the 12 go to 6?

OpenStudy (anonymous):

the x^12 = x^6 times x ^6 or (x^6)^2. so when x^12 is taken under sqrt, it becomes x^6